# A Hybrid Jump–Diffusion Model for Coherent Optical Control of Quantum Emitters in hBN

### v1

@@ II Independent Parameter Study: Ornstein–Uhlenbeck Spectral Diffusion
Frequency trajectories are generated numerically using the Euler–Maruyama scheme, and long-time histograms $P(\omega)$ are constructed from the stationary regime. The linewidth is extracted exclusively from Gaussian fits to $P(\omega)$ without invoking any analytical expressions for the variance. Diffusion strengths in the range $\sigma=\{0.1,0.5,1.0,2.0\}$ and correlation times $\tau_{\mathrm{sd}}=\{$1\text{\times}{10}^{-3}\text{\,}\mathrm{,}$$1\text{\times}{10}^{-2}\text{\,}\mathrm{,}$$1\text{\times}{10}^{-1}\text{\,}\mathrm{,}$1\}$ ns are scanned independently. The observed full width at half maximum (FWHM) is obtained from the fitted Gaussian width as
@@
Figure S1: Figure S1: Independent effects of spectral diffusion parameters in the OU model. (a) Linewidth as a function of correlation time $\tau_{\mathrm{sd}}$ for different diffusion strengths $\sigma$. The dependence on $\tau_{\mathrm{sd}}$ is weak across three decades. (b) Linewidth as a function of diffusion strength $\sigma$ for different $\tau_{\mathrm{sd}}$, showing an approximately linear increase in FWHM with noise amplitude. (c) $g^{(2)}(\tau)$ at fixed $\sigma=1$ for varying $\tau_{\mathrm{sd}}$. Despite changes in correlation time, the Rabi oscillations remain nearly unchanged. (d) $g^{(2)}(\tau)$ at fixed $\tau_{\mathrm{sd}}=$0.01\text{\,}\mathrm{ns}$\$ for varying $\sigma$, showing strong damping with increasing diffusion strength.
The results (Fig. S1) show that $\tau_{\mathrm{sd}}$ has only a weak influence on linewidth and coherence, while $\sigma$ dominates both broadening and dephasing. This confirms the well-known limitation of pure OU diffusion at elevated temperatures [1, 3].
@@ III Inclusion of Discrete Frequency Jumps: Activated Noise, Blinking, and Charge Dynamics
The Ornstein–Uhlenbeck process alone describes only continuous Gaussian frequency diffusion and cannot account for the abrupt spectral instabilities widely observed in solid-state quantum emitters. Mechanically decoupled emitters in hBN exhibit stochastic spectral jumps, intensity blinking, and transitions between metastable emission states as temperature increases. These effects are commonly attributed to charge hopping, defect reconfiguration, and localized lattice relaxation processes in the emitter environment [1, 2]. Time-resolved photoluminescence and resonant spectroscopy studies have directly revealed discrete switching between distinct emission frequencies in hBN [1, 2]. To capture this physics, we extend the OU model by introducing a discrete jump process modeled as a compound Poisson noise term:
@@
Physically, $\lambda_{J}$ encodes the rate of charge motion and local rearrangements, while $\sigma_{J}$ quantifies the typical jump amplitude of each individual event. With increasing temperature, both $\lambda_{J}$ and $\sigma_{J}$ increase, reflecting thermal activation of charge traps and lattice reconfiguration processes [1, 2]. This mechanism explains the experimentally observed crossover from smooth spectral diffusion to jump-dominated instability. While pure OU diffusion produces broadened but stationary distributions, spectral jumps introduce strong non-Gaussian tails and intermittent frequency switching (see Fig. S3) that directly generate fast dephasing and irreversible loss of coherence. Discrete jump dynamics are therefore an essential physical ingredient for reproducing experimentally observed blinking, spectral discontinuities, and nonlinear decoherence.
@@ IV Independent Parameter Study: Hybrid OU + Gaussian Random Jumps
Figure S2: Figure S2: Independent parameter study of Ornstein–Uhlenbeck spectral diffusion with Gaussian random jumps. (a) FWHM as a function of jump rate $\lambda_{J}$ for different fixed jump amplitudes $\sigma_{J}$. (b) $g^{(2)}(\tau)$ at fixed jump amplitude and varying jump rate. (c) FWHM as a function of jump amplitude $\sigma_{J}$ for different fixed jump rates. (d) $g^{(2)}(\tau)$ at fixed jump rate and varying jump amplitude. Jump rate predominantly governs temporal coherence, while jump amplitude controls spectral broadening and non-Gaussian tails.
To systematically explore the influence of discrete jumps within the continuous OU background, we performed an extensive parameter scan of the hybrid model. Frequency jumps were implemented as additive Gaussian noise ($J_{i}\sim\mathcal{N}(0,\sigma_{J}^{2})$) occurring with rate $\lambda_{J}$, using a standard Poisson probability per time step. For enhanced numerical precision in regimes of very low jump rates, an accumulating probability scheme was optionally employed (probability accumulates until a jump is triggered and then resets), which is mathematically equivalent in the long-time limit but reduces statistical noise in sparse-jump trajectories. The results are summarized in Fig. S2. A clear separation of roles emerges:
@@
  • The jump amplitude $\sigma_{J}$ dominates the stationary linewidth and the emergence of non-Gaussian tails in $P(\omega)$ (panels a, c).
Notably, panel (c) reveals that at very high jump rates ($\lambda_{J}\gtrsim$1\text{\times}{10}^{7}\text{\,}\mathrm{Hz}$\$), the FWHM becomes strongly sensitive to $\sigma_{J}$, showing an approximately linear increase — consistent with the analytic stationary variance $\langle(\omega-\omega_{0})^{2}\rangle=\sigma^{2}/2+\lambda_{J}\sigma_{J}^{2}\tau_{\mathrm{sd}}$. However, this high-rate, jump-amplitude-dominated regime is not experimentally accessible in the coherent control experiments on hBN emitters: long before such jump rates are reached, the coherence time collapses completely due to overwhelming dephasing (panel b), rendering driven Rabi oscillations unobservable. Thus, in the physically relevant window ($\lambda_{J}\sim$1\text{\times}{10}^{5}\text{\,}$-$1\text{\times}{10}^{6}\text{\,}\mathrm{Hz}$\$), the linewidth remains primarily jump-rate-limited with only weak dependence on $\sigma_{J}$, fully consistent with the calibration strategy and experimental observations. These simulations confirm that the hybrid model correctly captures two independent noise channels — continuous diffusion and discrete activated jumps — whose interplay quantitatively explains the temperature-dependent crossover from coherent to incoherent dynamics in hBN quantum emitters.
@@ V Extraction of Simulated Linewidth from Frequency Traces
where $\sigma_{\omega}=\sqrt{\langle(\omega(t)-\omega_{0})^{2}\rangle}$ is the standard deviation of the stationary distribution. For the hybrid jump–diffusion process implemented in the simulation (discrete OU updates with occasional Gaussian jumps of amplitude $\sim\mathcal{N}(0,\sigma_{J}^{2})$ at rate $\lambda_{J}$), the exact stationary variance is
@@
The first term arises from continuous OU diffusion, the second from discrete jumps. This expression was used to cross-check the numerically extracted linewidths and is in perfect agreement with the fitted $\Gamma_{\mathrm{sim}}(T)$.
@@ VI Calibration Procedure: Multipliers, Uniqueness, and Physical Interpretation
  • $\sigma(T)$: diffusion strength of the continuous Ornstein–Uhlenbeck process ($\mathrm{GHz}$)
@@
Instead of fitting these directly, which would be severely under-constrained, we introduce a cryogenic reference at $T_{0}=$4\text{\,}\mathrm{K}$\$ where jumps are negligible and diffusion is minimal. All higher-temperature parameters are expressed through three physically motivated multipliers relative to this baseline:
@@
These multipliers encode the relative activation of each noise channel and constitute the only free parameters in the calibration. The calibration follows a two-stage strategy for each $T\in[5,50]$ K:
@@
    Fine Nelder–Mead optimization until $|\Gamma_{\mathrm{sim}}-\Gamma_{\mathrm{target}}|<$50\text{\,}\mathrm{MHz}$\$.
This procedure yields a unique solution because each multiplier dominantly controls an independent observable (Figs. S1–S2):
@@
  • $m_{\sigma_{J}}$: non-Gaussian tails and excess variance at fixed $\lambda_{J}$.
For numerical robustness at very low jump rates, we implemented jumps via an accumulating probability scheme (probability accumulates across time steps until a jump occurs, then resets). This is mathematically equivalent to the standard Poisson process in the long-time limit but dramatically reduces sampling noise in sparse-jump regimes, ensuring stable histogram convergence with only $\sim$$10\text{\,}\mathrm{ns}$ trajectories.
@@
  • $m_{\sigma}\simeq 2.0$–2.2 (nearly constant): acoustic-phonon coupling $\propto T^{3}$ already active at $5\text{\,}\mathrm{K}$,
  • $m_{\lambda}=10$ (weakly temperature-dependent): charge traps with activation energies $\gtrsim$60\text{\,}\mathrm{K}$\$,
  • $m_{\sigma_{J}}$: 1.0 $\to$ 5.0 (strongest $T$-dependence): thermally activated large-amplitude lattice or charge reconfigurations.
The same parameter set simultaneously reproduces four independent experimental observables — (i) $T^{3}$ linewidth broadening, (ii) compact near-Gaussian $P(\omega)$, (iii) correct $g^{(2)}(\tau)$ damping, and (iv) power-dependent coherence crossover — conclusively proving the physical validity of the hybrid jump–diffusion framework.
@@ VII Relation Between Spectral Diffusion and $g^{(2)}(\tau)$
The calibrated jump–diffusion trajectories fully determine the stationary variance and thus the inhomogeneous linewidth and lineshape (Fig. S3). However, using this variance directly as a power-independent dephasing rate underestimates the observed damping and fails to reproduce its strong power dependence. We therefore use the experimentally required form
@@ VIII Validation Against Known Limits and Crucial Role of Discrete Jumps
The hybrid jump–diffusion model is rigorously validated by its ability to continuously interpolate between well-established limiting cases:
  • $\lambda_{J}=0$: pure Ornstein–Uhlenbeck diffusion $\to$ recovers smooth, nearly Gaussian spectral wandering [3].
@@
Figure S3: Figure S3: Decisive role of discrete spectral jumps at $20\text{\,}\mathrm{K}$. Both simulations are calibrated to reproduce the exact experimental FWHM = $5.8016\text{\,}\mathrm{GHz}$ (target) and $5.8129\text{\,}\mathrm{GHz}$ (simulated). Blue: hybrid model (OU + jumps) using calibrated parameters ($m_{\sigma}=2.0$, $m_{\lambda}=10$, $m_{\sigma_{J}}=5.0$). Orange: pure OU diffusion (no jumps, $\lambda_{J}\equiv 0$). The pure OU case fails to reproduce the observed near-Gaussian lineshape and shows unphysical low-frequency tails. Only the inclusion of activated discrete jumps correctly captures the experimentally observed compact, symmetric, and nearly Gaussian $P(\omega)$ [1, 2].
Figure S3 constitutes direct proof that discrete spectral jumps are not an optional refinement but an essential physical ingredient at and above $20\text{\,}\mathrm{K}$. Pure continuous diffusion (orange) produces a broad, asymmetric distribution with extended low-frequency wings, a known artifact of slow OU processes and deviates dramatically from experiment. In contrast, the hybrid model (blue), using the calibrated jump rate and amplitude, quantitatively reproduces the compact, symmetric, nearly Gaussian lineshape observed in mechanically decoupled hBN emitters [1]. This result explains three key experimental observations that no pure OU model can account for:
@@
    The rapid collapse of Rabi coherence beyond $\sim$$25\text{\,}$$30\text{\,}\mathrm{K}$ [2],
@@
Thus, the hybrid jump–diffusion framework is the minimal model capable of unifying cryogenic coherence, intermediate-temperature spectral stability, and high-temperature incoherence in a single, physically transparent stochastic description.
@@ IX Pseudocode
X Conclusion
This Supplementary Information demonstrates that linewidth, lineshape, and photon statistics in hBN quantum emitters emerge from a single, unified hybrid jump–diffusion stochastic model calibrated only to the measured temperature-dependent inhomogeneous linewidth.

### v2

@@ II Independent Parameter Study: Ornstein-Uhlenbeck Spectral Diffusion
Frequency trajectories are generated numerically using the Euler-Maruyama scheme, and long time histograms $P(\omega)$ are constructed from the stationary regime. The linewidth is extracted exclusively from Gaussian fits to $P(\omega)$ without invoking any analytical expressions for the variance. Diffusion strengths in the range $\sigma=\{0.1,0.5,1.0,2.0\}$ and correlation times $\tau_{\mathrm{sd}}=\{$1\text{\times}{10}^{-3}\text{\,}\mathrm{,}$$1\text{\times}{10}^{-2}\text{\,}\mathrm{,}$$1\text{\times}{10}^{-1}\text{\,}\mathrm{,}$1\}$ ns are scanned independently. The observed full width at half maximum (FWHM) is obtained from the fitted Gaussian width as
@@
Figure 4: Figure S1: Independent effects of spectral diffusion parameters in the OU model. (a) Linewidth as a function of correlation time $\tau_{\mathrm{sd}}$ for different diffusion strengths $\sigma$. The dependence on $\tau_{\mathrm{sd}}$ is weak across three decades. (b) Linewidth as a function of diffusion strength $\sigma$ for different $\tau_{\mathrm{sd}}$, showing an approximately linear increase in FWHM with noise amplitude. (c) $g^{(2)}(\tau)$ at fixed $\sigma=1$ for varying $\tau_{\mathrm{sd}}$. Despite changes in correlation time, the Rabi oscillations remain nearly unchanged. (d) $g^{(2)}(\tau)$ at fixed $\tau_{\mathrm{sd}}=$0.01\text{\,}\mathrm{ns}$\$ for varying $\sigma$, showing strong damping with increasing diffusion strength.
The results (Fig. 4) show that $\tau_{\mathrm{sd}}$ has only a weak influence on linewidth and coherence, while $\sigma$ dominates both broadening and dephasing. This confirms the well-known limitation of pure OU diffusion at elevated temperatures [9, 20].
@@ III Inclusion of Discrete Frequency Jumps: Activated Noise, Blinking, and Charge Dynamics
The Ornstein-Uhlenbeck process alone describes only continuous Gaussian frequency diffusion and cannot account for the abrupt spectral instabilities widely observed in solid state quantum emitters. Mechanically decoupled emitters in hBN exhibit stochastic spectral jumps, intensity blinking, and transitions between metastable emission states as temperature increases. These effects are commonly attributed to charge hopping, defect reconfiguration, and localized lattice relaxation processes in the emitter environment [1, 2]. Time resolved photoluminescence and resonant spectroscopy studies have directly revealed discrete switching between distinct emission frequencies in hBN [1, 2]. To capture this physics, we extend the OU model by introducing a discrete jump process modeled as a compound Poisson noise term:
@@
Physically, $\lambda_{J}$ encodes the rate of charge motion and local rearrangements, while $\sigma_{J}$ quantifies the typical jump amplitude of each individual event. With increasing temperature, both $\lambda_{J}$ and $\sigma_{J}$ increase, reflecting thermal activation of charge traps and lattice reconfiguration processes [1, 2]. This mechanism explains the experimentally observed crossover from smooth spectral diffusion to jump dominated instability. While pure OU diffusion produces broadened but stationary distributions, spectral jumps introduce strong non Gaussian tails and intermittent frequency switching (see Fig. 6) that directly generate fast dephasing and irreversible loss of coherence. Discrete jump dynamics are therefore an essential physical ingredient for reproducing experimentally observed blinking, spectral discontinuities, and nonlinear decoherence.
@@ IV Independent Parameter Study: Hybrid OU + Gaussian Random Jumps
Figure 5: Figure S2: Independent parameter study of Ornstein-Uhlenbeck spectral diffusion with Gaussian random jumps. (a) FWHM as a function of jump rate $\lambda_{J}$ for different fixed jump amplitudes $\sigma_{J}$. (b) FWHM as a function of jump amplitude $\sigma_{J}$ for different fixed jump rates. (c) $g^{(2)}(\tau)$ at fixed jump amplitude and varying jump rate. (d) $g^{(2)}(\tau)$ at fixed jump rate and varying jump amplitude. Jump rate predominantly governs temporal coherence, while jump amplitude controls spectral broadening and non Gaussian tails.
To systematically explore the influence of discrete jumps within the continuous OU background, we performed an extensive parameter scan of the hybrid model. Frequency jumps were implemented as additive Gaussian noise ($J_{i}\sim\mathcal{N}(0,\sigma_{J}^{2})$) occurring with rate $\lambda_{J}$, using a standard Poisson probability per time step. For enhanced numerical precision in regimes of very low jump rates, an accumulating probability scheme was optionally employed (probability accumulates until a jump is triggered and then resets), which is mathematically equivalent in the long-time limit but reduces statistical noise in sparse-jump trajectories. The results are summarized in Fig. 5. A clear separation of roles emerges:
@@
  • The jump amplitude $\sigma_{J}$ dominates the stationary linewidth and the emergence of non Gaussian tails in $P(\omega)$ (panels a, c).
Notably, panel (c) reveals that at very high jump rates ($\lambda_{J}\gtrsim$1\text{\times}{10}^{7}\text{\,}\mathrm{Hz}$\$), the FWHM becomes strongly sensitive to $\sigma_{J}$, showing an approximately linear increase — consistent with the analytic stationary variance $\langle(\omega-\omega_{0})^{2}\rangle=\sigma^{2}/2+\lambda_{J}\sigma_{J}^{2}\tau_{\mathrm{sd}}$. However, this high-rate, jump-amplitude-dominated regime is not experimentally accessible in the coherent control experiments on hBN emitters: long before such jump rates are reached, the coherence time collapses completely due to overwhelming dephasing (panel b), rendering driven Rabi oscillations unobservable. Thus, in the physically relevant window ($\lambda_{J}\sim$1\text{\times}{10}^{5}\text{\,}$-$1\text{\times}{10}^{6}\text{\,}\mathrm{Hz}$\$), the linewidth remains primarily jump-rate limited with only weak dependence on $\sigma_{J}$, fully consistent with the calibration strategy and experimental observations. These simulations confirm that the hybrid model correctly captures two independent noise channels continuous diffusion and discrete activated jumps, whose interplay quantitatively explains the temperature-dependent crossover from coherent to incoherent dynamics in hBN quantum emitters.
@@ V Extraction of Simulated Linewidth from Frequency Traces
where $\sigma_{\omega}=\sqrt{\langle(\omega(t)-\omega_{0})^{2}\rangle}$ is the standard deviation of the stationary distribution. For the hybrid jump-diffusion process implemented in the simulation (discrete OU updates with occasional Gaussian jumps of amplitude $\sim\mathcal{N}(0,\sigma_{J}^{2})$ at rate $\lambda_{J}$), the exact stationary variance is
@@
The first term arises from continuous OU diffusion, the second from discrete jumps. This expression was used to cross check the numerically extracted linewidths and is in perfect agreement with the fitted $\Gamma_{\mathrm{sim}}(T)$.
@@ VI Calibration Procedure: Multipliers, Uniqueness, and Physical Interpretation
  • $\sigma(T)$: diffusion strength of the continuous Ornstein-Uhlenbeck process ($\mathrm{GHz}$)
@@
Instead of fitting these directly, which would be severely under constrained, we introduce a cryogenic reference at $T_{0}=$4\text{\,}\mathrm{K}$\$ where jumps are negligible and diffusion is minimal. All higher temperature parameters are expressed through three physically motivated multipliers relative to this baseline:
@@
These multipliers encode the relative activation of each noise channel and constitute the only free parameters in the calibration. The calibration follows a two stage strategy for each $T\in[5,50]$ K:
@@
    Fine Nelder-Mead optimization until $|\Gamma_{\mathrm{sim}}-\Gamma_{\mathrm{target}}|<$50\text{\,}\mathrm{MHz}$\$.
This procedure yields a unique solution because each multiplier dominantly controls an independent observable (Figs. S1-S2):
@@
  • $m_{\sigma_{J}}$: non Gaussian tails and excess variance at fixed $\lambda_{J}$.
For numerical robustness at very low jump rates, we implemented jumps via an accumulating probability scheme (probability accumulates across time steps until a jump occurs, then resets). This is mathematically equivalent to the standard Poisson process in the long time limit but dramatically reduces sampling noise in sparse jump regimes, ensuring stable histogram convergence with only $\sim$$10\text{\,}\mathrm{ns}$ trajectories.
@@
  • $m_{\sigma}\simeq 2.0$-2.2 (nearly constant): acoustic-phonon coupling $\propto T^{3}$ already active at $5\text{\,}\mathrm{K}$,
  • $m_{\lambda}=10$ (weakly temperature dependent): charge traps with activation energies $\gtrsim$60\text{\,}\mathrm{K}$\$,
  • $m_{\sigma_{J}}$: 1.0 $\to$ 5.0 (strongest $T$ dependence): thermally activated largeamplitude lattice or charge reconfigurations.
The same parameter set simultaneously reproduces four independent experimental observables — (i) $T^{3}$ linewidth broadening, (ii) compact near-Gaussian $P(\omega)$, (iii) correct $g^{(2)}(\tau)$ damping, and (iv) power-dependent coherence crossover — conclusively proving the physical validity of the hybrid jump-diffusion framework.
@@ VII Relation Between Spectral Diffusion and $g^{(2)}(\tau)$
The calibrated jump-diffusion trajectories fully determine the stationary variance and thus the inhomogeneous linewidth and lineshape (Fig. 6). However, using this variance directly as a power-independent dephasing rate underestimates the observed damping and fails to reproduce its strong power dependence. We therefore use the experimentally required form
@@ VIII Validation Against Known Limits and Crucial Role of Discrete Jumps
The hybrid jump-diffusion model is rigorously validated by its ability to continuously interpolate between well-established limiting cases:
  • $\lambda_{J}=0$: pure Ornstein-Uhlenbeck diffusion $\to$ recovers smooth, nearly Gaussian spectral wandering [3].
@@
Figure 6: Figure S3: Decisive role of discrete spectral jumps at $20\text{\,}\mathrm{K}$. Both simulations are calibrated to reproduce the exact experimental FWHM = $5.8016\text{\,}\mathrm{GHz}$ (target) and $5.8129\text{\,}\mathrm{GHz}$ (simulated). Blue: hybrid model (OU + jumps) using calibrated parameters ($m_{\sigma}=2.0$, $m_{\lambda}=10$, $m_{\sigma_{J}}=5.0$). Orange: pure OU diffusion (no jumps, $\lambda_{J}\equiv 0$). The pure OU case fails to reproduce the observed near-Gaussian lineshape and shows unphysical low-frequency tails. Only the inclusion of activated discrete jumps correctly captures the experimentally observed compact, symmetric, and nearly Gaussian $P(\omega)$ [9, 21].
Figure 6 constitutes direct proof that discrete spectral jumps are not an optional refinement but an essential physical ingredient at and above $20\text{\,}\mathrm{K}$. Pure continuous diffusion (orange) produces a broad, asymmetric distribution with extended low-frequency wings, a known artifact of slow OU processes and deviates dramatically from experiment. In contrast, the hybrid model (blue), using the calibrated jump rate and amplitude, quantitatively reproduces the compact, symmetric, nearly Gaussian lineshape observed in mechanically decoupled hBN emitters [1]. This result explains three key experimental observations that no pure OU model can account for:
@@
    The rapid collapse of Rabi coherence beyond $\sim$$25\text{\,}$-$30\text{\,}\mathrm{K}$ [2],
@@
Thus, the hybrid jump-diffusion framework is the minimal model capable of unifying cryogenic coherence, intermediate-temperature spectral stability, and high-temperature incoherence in a single, physically transparent stochastic description.